# Simulating Persuasive Dialogues on Meat Reduction with Generative Agents

### v1

@@ 1 Introduction
Reducing meat consumption offers substantial benefits for human and planetary health, yet entrenched social norms continue to place meat at the center of shared meals (Godfray et al. 2018). Vegans, vegetarians, and flexitarians have the potential to challenge these norms and drive important social change (Judge et al. 2024). However, in mixed-diet social settings, they often self-silence to avoid social costs (e.g., Bolderdijk and Cornelissen 2022; Romo and Donovan-Kicken 2012). To date, the development of communication strategies that promote meat reduction while minimizing social costs has required human participants at each stage of the process (e.g., Carfora et al. 2019; Pabian et al. 2020). This is not only very costly—both financially for the researcher and in terms of participant effort—but also severely restricts the range of communication strategies and participant diversity that can be explored.
@@
We present work in progress on simulating multi-round dialogues on meat reduction between generative agents based on large language models (LLMs), as shown in Figure 1. Our primary focus in this paper lies in validating the simulation of the generative agents internally and against existing data from previous experiments with human participants.
@@ Results
Initial experiments with the Llama 3 family of models indicate that generative agents can effectively simulate persuasive communication, producing reliable and valid response patterns across key psychological constructs. However, issues such as uniformity in some constructs and the potential influence of instruction-tuning on strategies will require further investigation. Although the approach is promising, it will be necessary to refine the simulation setup to eventually make meaningful statements about human behavior. At the same time, generative agent-based models can be a promising tool for identifying novel communication strategies for meat reduction, tailored to highly specific participant groups, to then be tested in subsequent studies with human participants.
@@ 2 Background
Persuasion for Meat Reduction
The Theory of Planned Behavior (TPB; Ajzen 1991) explains behavior through three core constructs: Attitudes, Subjective Norms, and Perceived Behavioral Control. These shape behavioral Intentions, which in turn predict actual behavior. On average, intentions account for 28% of variance in health behavior (Sheeran et al. 2005). In meat reduction, these variables reflect recipient traits as follows: attitudes map personal beliefs about meat, norms relate to perceived expectations of others, and control captures confidence in one’s ability to change eating habits. From theory and previous research, all three constructs reliably predict the intention to eat less meat, which in turn predicts actual meat consumption (Berndsen and Van Der Pligt 2005; Saba and Di Natale 1998). We expect individuals with favorable attitudes, strong normative support, and high perceived control to be more receptive to meat reduction messages. We define the TPB as our theoretical framework and the corresponding variables as our outcomes for persuasion. Consequently, we use validated instruments to assess these constructs reliably. As part of our simulation, we extract survey responses from the agents to be persuaded (Recipient), see Appendix A for the complete questionnaire.
@@
Social Simulations with Generative Agents
A rapidly growing body of research investigates the feasibility of using large language models (LLMs) to model survey responses (e.g., Argyle et al. 2023; Bisbee et al. 2023; Ahnert et al. 2025), or the outcomes of experimental studies (e.g., Hewitt et al. 2024; Aher, Arriaga, and Kalai 2024). LLMs promise to facilitate novel exploratory studies of human behavior (Bail 2024), but recent research indicates that LLMs might not always accurately mimic human study participants (Tjuatja et al. 2024), and, for instance, “flatten” the reported attitudes of identity groups (Wang, Morgenstern, and Dickerson 2024).
@@ 3 Approach
We simulate meat reduction dialogues between two generative agents (Persuader and Recipient) as shown in Figure 2. Every conversation is started by the Persuader and consists of 5 rounds, in which both agents produce 1 message each. To produce a new message, an agent goes through the following steps: it (1) performs internal reflection; Recipient agents (2) answer a questionnaire; and finally, an agent (3) generates a response to the other agent. The Persuader has explicit instructions to persuade the other agent to reduce meat, while the Recipient is instructed to adopt one of the personas shown in Table 1. We explain the simulation process in the following and publish all respective prompts as part of our code under MIT license. [^3] https://github.com/dess-mannheim/MeatlessAgents
@@
To perform internal reflection, all agents receive a transcript of the conversation so far and are instructed to reflect on which emotions were evoked, which core values were challenged or supported, and which questions or uncertainties there are. The questionnaire is only filled out by the Recipients and consists of questions that have previously been used with human study participants. It measures our 6 central constructs: Attitudes, Subjective Norms, and Behavioral Control of the Recipient; the Recipient’s Intention to Reduce Meat Consumption; as well as the Social Closeness and the Social Attractiveness of the Persuader, as perceived by the Recipient. The first 4 constructs comprise the aspects of the TPB that we can measure in our simulation, while the latter 2 constructs measure Social Costs (see Appendix A for the full questionnaire). Finally, each agent generates a response to the other agent. This response takes into account only the agent’s persona, the transcript of the conversation so far, and the agent’s internal reflection. To ensure that the questionnaire does not affect the simulation, we let agents “forget” about it.
@@
For our initial analyzes, we simulate 200 conversations for both contrasting Recipient personas that are shown in Table 1, i.e., 400 conversations in total. We extract 10 independent questionnaire answers from the Recipient persona to improve robustness. We opt for open-weight LLMs to facilitate replication (Barrie, Palmer, and Spirling 2024), and run our simulation with 3 distinct LLMs to investigate the impact of model size: Llama 3.3 70B, Llama 3.1 8B, and Llama 3.2 3B. For the questionnaire, we employ structured outputs tailored to the respective answer options, to ensure that we obtain valid responses even from the smallest model. We run all simulations with Llama’s default temperature of $0.6$ and use vllm with automatic prefix caching to improve model throughput—the combined runtime for all simulations on two NVIDIA H100 in parallel was $\approx 70$ hours.
@@ 4 Results
[…]
@@
Comparing Simulated and Human Data
To assess the similarity between simulated and human data, we compared the mean values of key constructs with those reported in previous studies using the same measures with human participants. As we did not have access to the original datasets, the comparisons were based on published mean values. A scatterplot of the simulated versus reported means is shown in Figure 3. Most constructs produced higher mean values in the simulation compared to human data, with the exception of Attitudes, where the simulated mean was slightly lower. The largest discrepancy was observed for Subjective Norms, which were generally low in both cases, but especially in the human samples. In general, the points cluster closely around the identity line, indicating a high degree of similarity between the simulated and human-generated data, particularly in the relative positioning of the constructs. This supports the validity of the simulation in capturing psychologically plausible response patterns. However, these findings should be interpreted as preliminary evidence.
@@
[…]
Figure 5: Recipients’ Intention to Reduce Meat. Combined survey responses after each conversation round, where $\Diamond$ indicates the mean over all conversations per round. Over the first rounds of the conversation, the Hard to Persuade Recipient shows an initially decreasing intention to reduce meat, similar to human participants (Shen, Sheer, and Li 2015).
@@
Figure 5 shows the reported intention to reduce meat consumption—our main outcome—and each point represents the mean response of a simulated Recipient across the 3 respective questionnaire items and 10 repeated survey responses. As expected, the Hard to Persuade Recipient indicates a much smaller intention for meat reduction than the Easy to Persuade Participant. For Llama 3.3 70B, we also observe much more variance in the responses from the Hard to Persuade Recipient. Surprisingly, during the first rounds of the conversation, the simulated survey responses for the Hard to Persuade Recipient show a decreasing intention to reduce meat consumption. We also observe this trend, although to a lesser extent, for the smaller Llama models (see Appendix Figure 9). This could be interpreted as a backlash to the persuasion attempt, a pattern that is also observed with human participants (Shen, Sheer, and Li 2015). After approximately 3 conversation rounds, we observe an increasing mean intention for meat reduction, across both participants and all LLMs. This trend continues in round 5 for the Hard to Persuade Recipient and Llama 3.3 70B, which we take as a reason to extend future simulation to more rounds.
@@ 5 Conclusion
Our preliminary findings demonstrate that generative agents can engage in plausible discussions about meat reduction. However, limitations such as uniformity in subjective norm responses require further investigation. As next steps, we will extend the simulation to more rounds and run targeted simulations in which we instruct the Persuader to employ a specific persuasion strategy.
@@ Appendix A Questionnaires
Figure 6: Theory of Planned Behavior and Social Costs. Behavior itself (i.e., meat reduction) is unobservable to us in our simulation, but previous research has established that self-reported intention to reduce meat consumption adequately predicts actual reduction (Berndsen and Van Der Pligt 2005; Saba and Di Natale 1998).
@@
Social attractiveness was measured with seven items adapted from Thürmer and McCrea (2023), using a 7-point Likert scale:
@@ Appendix B Additional Results
Scale #items Llama 3.3 70B Llama 3.1 8B Llama 3.2 3B Human Samples Reported by
@@
Social Attractiveness 7 .98 .93 .96 .91 Thürmer and McCrea (2023)
@@
[…]

### v2

@@ 1 Introduction
Reducing meat consumption offers substantial benefits for human and planetary health, yet entrenched social norms continue to place meat at the center of shared meals (Godfray et al. 2018). Vegans, vegetarians, and flexitarians have the potential to challenge these norms and drive important social change (Judge et al. 2024). However, in mixed-diet social settings, they often self-silence to avoid social costs (e.g., Bolderdijk and Cornelissen 2022; Romo and Donovan-Kicken 2012). To date, the development of communication strategies that promote meat reduction while minimizing social costs has required human participants at each stage of the process (e.g., Carfora et al. 2019; Pabian et al. 2020). This is not only very costly—both financially for the researcher and in terms of participant effort—but also severely restricts the range of communication strategies and participant diversity that can be explored. Generative Agent simulations based on large language models (LLMs) have the potential to overcome these limitations (Bail 2024).
@@
We present work in progress on simulating multi-round dialogues on meat reduction between generative agents based on LLMs, as shown in Figure 1. Related work on generative agent simulations already shows promising results (Park et al. 2022; Törnberg et al. 2023). Still, it remains an open question how well such simulations can inform subsequent studies with human participants, especially in our context of dialogues on meat reduction. This paper therefore addresses the following Research Question:
To what extent can generative agent simulations reliably reproduce persuasive human dialogues about meat reduction? To answer this, we validate the behavior of our agents both internally and against data from prior experiments with human participants. This validation is a necessary step before using these simulations to inform the design of future studies involving human subjects.
@@ Results
Initial experiments with the Llama 3 family of models indicate that generative agents can effectively simulate persuasive communication, producing reliable and valid response patterns across key psychological constructs. However, issues such as uniformity in some constructs and the potential influence of instruction-tuning on strategies will require further investigation.
Generative agent-based models can be a promising tool for identifying novel communication strategies for meat reduction, tailored to highly specific participant groups, to then be tested in subsequent studies with human participants.
@@ 2 Background
Figure 2: Constructs Measured for the Recipient, Comprising Theory of Planned Behavior (TPB) and Social Costs. Behavior itself (i.e., meat reduction) is unobservable to us in our simulation, but previous research has established that self-reported Intention to Reduce Meat—our main outcome—adequately predicts actual reduction (Berndsen and Van Der Pligt 2005; Saba and Di Natale 1998).
Studies on Meat Reduction
The Theory of Planned Behavior (TPB; Ajzen 1991) explains behavior through three core constructs: Attitudes, Subjective Norms, and Perceived Behavioral Control. These shape behavioral Intentions, which in turn predict actual behavior. On average, intentions account for 28% of variance in health behavior (Sheeran et al. 2005). In meat reduction, these variables reflect recipient traits as follows: attitudes map personal beliefs about meat, norms relate to perceived expectations of others, and control captures confidence in one’s ability to change eating habits. From theory and previous research, all three constructs reliably predict the intention to eat less meat, which in turn predicts actual meat consumption (Berndsen and Van Der Pligt 2005; Saba and Di Natale 1998).
Social Simulations with Generative Agents
A rapidly growing body of research investigates the feasibility of using large language models (LLMs) to model survey responses (e.g., Argyle et al. 2023; Bisbee et al. 2023; Ahnert et al. 2025), or the outcomes of experimental studies (e.g., Hewitt et al. 2024; Aher, Arriaga, and Kalai 2024). LLMs promise to facilitate novel exploratory studies of human behavior (Bail 2024), but recent research indicates that LLMs might not always accurately mimic human study participants (Tjuatja et al. 2024). It remains an open question how well socio-demographically prompted LLMs can represent human study participants (Sen et al. 2025)—in particular, LLMs were shown to “flatten” the reported attitudes of identity groups (Wang, Morgenstern, and Dickerson 2024).
@@
3 Approach
Psychological Constructs
We define the Theory of Planned Behavior (TPB) as our theoretical framework and the corresponding variables as our outcomes for persuasion, as shown in Figure 2. Consequently, we use validated instruments to assess these constructs reliably. As part of our simulation, we extract survey responses from the agent to be persuaded (Recipient), with Intention to Reduce Meat as our main outcome. The questionnaire measures our 6 central constructs: Attitudes, Subjective Norms, and Behavioral Control of the Recipient; the Recipient’s Intention to Reduce Meat Consumption; as well as the Social Closeness and the Social Attractiveness of the Persuader, as perceived by the Recipient. The first 4 constructs comprise the aspects of the TPB that we can measure in our simulation, while the latter 2 constructs measure Social Costs—see Appendix A for the full questionnaire. We expect individuals with favorable attitudes, strong normative support, and high perceived control to be more receptive to meat reduction messages.
@@ 3 Approach
We simulate meat reduction dialogues between two generative agents (Persuader and Recipient) as shown in Figure 3. Every conversation is started by the Persuader and consists of 5 rounds, in which both agents produce 1 message each. To produce a new message, an agent goes through the following steps: it (1) performs internal reflection; Recipient agents (2) answer a questionnaire; and finally, an agent (3) generates a response to the other agent. The Persuader has explicit instructions to persuade the other agent to reduce meat, while the Recipient is instructed to adopt one of the personas shown in Table 1. In our initial simulations, we do not prompt the Persuader to follow a specific persuasion strategy. We explain the simulation process in the following and publish all respective prompts as part of our code under MIT license. [^2] https://github.com/dess-mannheim/MeatlessAgents
@@
To perform internal reflection, all agents receive a transcript of the conversation so far and are instructed to reflect on which emotions were evoked, which core values were challenged or supported, and which questions or uncertainties there are. Only the Recipient agent, then, fills out the questionnaire described above. Finally, each agent generates a response to the other agent. This response takes into account only the agent’s persona, the transcript of the conversation so far, and the agent’s internal reflection. To ensure that the questionnaire does not affect the simulation, we remove it from the LLM’s context.
@@
For our initial analyzes, we simulate 200 dialogues for both contrasting Recipient personas that are shown in Table 1, i.e., 400 conversations in total, with 10 messages exchanged in each dialogue. We extract 10 independent questionnaire answers from the Recipient persona to improve robustness. We opt for open-weight LLMs to facilitate replication (Barrie, Palmer, and Spirling 2024), and run our simulation with 3 distinct LLMs to investigate the impact of model size: Llama 3.3 70B, Llama 3.1 8B, and Llama 3.2 3B. For the questionnaire, we employ structured outputs tailored to the respective answer options, to ensure that we obtain valid responses even from the smallest model. We run all simulations with Llama’s default temperature of $0.6$ and use vllm with automatic prefix caching to improve model throughput—the combined runtime for all simulations on two NVIDIA H100 in parallel was $\approx 70$ hours.
Construct Original Study n
Attitude & Intention to Reduce Meat Pabian et al. (2020) 47
Subjective Norms & Behavioral Control Wyker and Davison (2010) 204
Social Attractiveness Thürmer, Stadler, and McCrea (2022) 260
Social Closeness Monin, Sawyer, and Marquez (2008) 70
Threat of Freedom Dillard and Shen (2005) 196
Meat Attachment Graça, Calheiros, and Oliveira (2015) 318
Table 2: Studies with Human Participants Used for External Validation. We compare published means against the survey responses from our generative agent simulation.
Validation
To assess whether generative agent simulations can meaningfully replicate persuasive human dialogues about meat reduction, we implement a two-step validation strategy. First, we perform an internal validation by evaluating whether the agents’ responses exhibit psychometric properties comparable to those observed in human data. Using established psychological constructs and measurement instruments, we assess internal consistency, convergent and discriminant validity, and distributional plausibility. Second, we conduct an external validation by comparing the simulated results against empirical data from prior human studies with similar persuasion tasks (see Table 2). This dual approach enables us to examine both the internal coherence and the empirical realism of the simulated dialogues. Crucially, this validation serves as a necessary foundation: only if our simulation proves sufficiently reliable and valid can it be used to inform the design of future experiments—with real human participants—under comparable conditions.
@@ 4 Results
To assess the internal consistency of the measured constructs, we calculated Cronbach’s Alpha coefficients for each scale. The results indicated excellent reliability for most constructs (e.g. Meat Attachment, Attitude, Intention, Social Attractiveness, and Social Closeness $\alpha s\geq.98$). In contrast, Subjective Norms ($\alpha=.76$) and Behavioral Control ($\alpha=.70$) showed comparatively lower but still acceptable internal consistency. A detailed overview of the reliability coefficients, including comparisons with the values obtained from human samples, is provided in Appendix Table 3.
Comparing Simulated and Human Data
To assess the similarity between simulated and human data, we compared the mean values of key constructs with those reported in previous studies using the same measures with human participants. As we did not have access to the original datasets, the comparisons were based on published mean values. Table 2 shows the original studies belonging to the respective constructs with their respective number of human participants. A scatterplot of the simulated versus reported means is shown in Figure 4. Most constructs produced higher mean values in the simulation compared to human data, with the exception of Attitudes, where the simulated mean was slightly lower. The largest discrepancy was observed for Subjective Norms, which were generally low in both cases, but especially in the human samples. In general, the points cluster closely around the identity line, indicating a high degree of similarity between the simulated and human-generated data, particularly in the relative positioning of the constructs. This supports the validity of the simulation in capturing psychologically plausible response patterns. However, these findings should be interpreted as preliminary evidence.
@@
To further assess the construct validity of the simulated responses, we examined the intercorrelations between all measured variables, as visualized in Figure 5. The results reveal strong positive correlations between constructs that are theoretically linked, such as Attitude, Intention, Social Attractiveness, and Social Closeness (e.g., $r=.96$ between Attitude and Intention). This supports convergent validity. Correlations with less strongly related constructs, such as relations with Subjective Norms, are notably lower (e.g., $r=.52$ between Behavioral Control and Subjective Norms), supporting discriminant validity. These two constructs were deliberately included as control variables to assess whether the central constructs form a conceptually distinct cluster. However, some correlations between central variables are exceptionally high (approaching or exceeding $r=.95$), which is uncommon in empirical data and may reflect the polarized nature of the simulated personas. This could lead to overly homogeneous response patterns, limiting generalizability, and possibly inflating convergence artificially. Future simulations could benefit from a broader and more nuanced range of personas to capture more natural variance.
@@
Figure 6 shows the reported intention to reduce meat consumption—our main outcome—and each point represents the mean response of a simulated Recipient across the 3 respective questionnaire items and 10 repeated survey responses. As expected, the Hard to Persuade Recipient indicates a much smaller intention for meat reduction than the Easy to Persuade Participant. For Llama 3.3 70B, we also observe much more variance in the responses from the Hard to Persuade Recipient. Surprisingly, during the first rounds of the conversation, the simulated survey responses for the Hard to Persuade Recipient show a decreasing intention to reduce meat consumption. This could be interpreted as a backlash to the persuasion attempt, a pattern that is also observed with human participants (Shen, Sheer, and Li 2015). After approximately 3 conversation rounds, we observe an increasing mean intention for meat reduction, across both participants and all LLMs. This trend continues in round 5 for the Hard to Persuade Recipient and Llama 3.3 70B, which we take as a reason to extend future simulation to more rounds.
Figure 6: Recipients’ Intention to Reduce Meat. Combined survey responses after each conversation round, where $\Diamond$ indicates the mean over all conversations per round. Following our expectations, the Easy to Persuade Recipient responds with a generally high intention to reduce meat. Over the first rounds of the conversation, the Hard to Persuade Recipient shows an initially decreasing intention to reduce meat, similar to human participants (Shen, Sheer, and Li 2015).
@@ 5 Conclusion
Our preliminary findings demonstrate that generative agents can engage in plausible discussions about meat reduction. However, limitations such as uniformity in subjective norm responses require further investigation. As next steps, we will extend the simulation to more rounds and run targeted simulations in which we instruct the Persuader to employ a specific persuasion strategy. Additionally, we will investigate more diverse personas.
@@ Appendix A Questionnaires
Social attractiveness was measured with seven items adapted from Thürmer, Stadler, and McCrea (2022), using a 7-point Likert scale:
@@ Appendix B Additional Results
Scale #items Llama 3.3 70B Llama 3.1 8B Llama 3.2 3B Human Samples Original Study
@@
Social Attractiveness 7 .98 .93 .96 .91 Thürmer, Stadler, and McCrea (2022)
@@
Table 3: Reliability of psychological constructs decreases with smaller Llama models. Cronbach’s Alpha values for each scale are compared across Llama models (3.3 70B, 3.1 8B, 3.2 3B) and human participant data to assess internal consistency. While larger models consistently show high reliability ($\alpha\geq.70$), several scales fall below the acceptable threshold in smaller models (shown in red)—most notably Behavioral Control, and Threat of Freedom. This suggests a degradation in the coherent representation of multi-item constructs as model size decreases.